OpenStudy (anonymous):

can you combine the radical 3√6-2√2

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

not like terms right?

OpenStudy (anonymous):

\[=\sqrt{2}(3\sqrt{3}-2)\]

OpenStudy (anonymous):

You can`t because its not the same number under the radical

OpenStudy (anonymous):

\[\sqrt{6}=\sqrt{3}*\sqrt{2}\]

OpenStudy (anonymous):

are we good?

OpenStudy (anonymous):

how is it that?

OpenStudy (anonymous):

\[you have \sqrt{6}=\sqrt{3*2}=\sqrt{3}*\sqrt{2}\] see it?

OpenStudy (anonymous):

How do you combine \[\sqrt{3}\] and \[\sqrt{2}\]?

OpenStudy (anonymous):

\[3\sqrt{6}-2\sqrt{2}=3\sqrt{3*2}-2\sqrt{2}=3\sqrt{3}\sqrt{2}-2\sqrt{2}=\sqrt{2}*(3\sqrt{3}-2)\]

OpenStudy (anonymous):

\[\sqrt{2}\] is common multiplier